An electric dipole with dipole moment P→ = (3.00î + 4.02ĵ )(1.48 × 10-30C·m) is in an electric field E→ = (4100 N/C)î. (a) What is the potential energy of the electric dipole? (b) What is the magnitude of torque acting on it? (c) If an external agent turns the dipole until its electric dipole moment is P→= (-4.02î + 3.00ĵ)(1.48 × 10-30 C·m), how much work is done by the agent?
Answer #1

Potential energy = P.E= (3i + 4.02 j ).4100i = - 12300J

Torque = P x E = (3i +4.02j )× 4100i= - 16482 J

Final Potential energy = P.E = - 16482 J

Initial potential energy = - 12300J

Work done = change in energy = -16482 + 12300= - 4182J

Work done by agent = 4182J